OpenStudy (hlpmeihlpu123):

Line A is represented by the equation given below: x + y = 3 What is most likely the equation for line B, so that the set of equations has infinitely many solutions? A. 3x + 3y = 3 B. 3x + y = 3 C. x + y = 9 D. 3x + 3y = 9

OpenStudy (hlpmeihlpu123):

i think its c

OpenStudy (hlpmeihlpu123):

hlp plzz will medal

OpenStudy (hlpmeihlpu123):

@ganeshie8 @Mehek14

OpenStudy (ijlal):

its D will show yo how :)

OpenStudy (hlpmeihlpu123):

thank you

OpenStudy (hlpmeihlpu123):

i have just one more will you help

OpenStudy (ijlal):

graph of x+y=3

OpenStudy (ijlal):

graph of 3x+3y=9

OpenStudy (hlpmeihlpu123):

A set of equations is given below: Equation C: y = 7x + 12 Equation D: y = 7x + 2 How many solutions are there to the given set of equations?

OpenStudy (hlpmeihlpu123):

thank you

OpenStudy (ijlal):

if you see both the graphs they are the same :) which means they both have infinetly many solutions as all the points will satisfy both the equations hope you understand :)

OpenStudy (hlpmeihlpu123):

is that the answer for the second one

OpenStudy (ijlal):

no its the answer for the first one i am solving the second one a moment pls

OpenStudy (hlpmeihlpu123):

ok srry

OpenStudy (ijlal):

well this also is the answer to the 2nd one also as both equations are same they both have infinitely many solutions :)

OpenStudy (hlpmeihlpu123):

thank you is that program that you made the graph with free

OpenStudy (ijlal):

yep it is

OpenStudy (hlpmeihlpu123):

what is it called

OpenStudy (ijlal):

download from this link

OpenStudy (hlpmeihlpu123):

thank you

OpenStudy (ijlal):

most welcome :)